Prove that ((x+y+z)/3)^(x+y+z) ≤ x^x y^y z^z ≤ ((x^2+y^2+z^2)/(x+y+z))^(x+y+z) Where x, y, z element of N
1
Expert's answer
2020-05-28T16:07:08-0400

To solve this problem, we will use

AM ≥GM twice.

Firstly, let us consider

x,x,. ... x-times, y,y,....... y-times,z,z.... z-times.

AM of above numbers is "\frac {x.x + y.y + z.z} {x+y+z}" = "\frac {x² + y² + z²} {x+y+z}"

GM of above numbers = ("\ x^x . y^y . z^z" )"\ ^ \frac {1} {(x+y+z)}"

Since AM ≥ GM,

"\frac {x²+y²+z²} {x+y+z}" ≥ ("\ x^x . y^y . z^z" )"\ ^ \frac {1} {(x+y+z)}"

=> ( "\frac {x²+y²+z²} {x+y+z}" )"\ ^ {x+y+z}""\ x^x . y^y . z^z" -----------> (1)

To prove  "(\frac{x+y+z}{3})^{x+y+z} \leq x^x y^y z^z", secondly, let us consider

1/x, 1/x,. ... x-times, 1/y, 1/y,....... y-times,1/z, 1/z.... z-times.

AM of above numbers is "\frac {x.(1/x) + y.(1/y) + z.(1/z)} {x+y+z}" = "\frac {3} {x+y+z}"

GM of above numbers =

"\ [ (1/x)^x . (1/y)^y . (1/z)^z ]" "\ ^ \frac {1} {(x+y+z)}"

= [ "\frac {1} {x^xy^yz^z}" ]"\ ^ \frac {1} {(x+y+z)}"

Since AM ≥ GM

"\frac {3} {x+y+z}" ≥ ["\frac {1} {x^xy^yz^z}" ]"\ ^ \frac {1} {(x+y+z)}"

=> ("\frac {3} {x+y+z}" )(x+y+z)"\frac {1} {x^xy^yz^z}"

Reciprocating,

"\ x^x . y^y . z^z" ≥ ("\frac {x+y+z} {3}" )(x+y+z) ---------> (2)

By inequality (1) and (2)

("\frac {x²+y²+z²} {x+y+z}" )(x+y+z) "\ x^x . y^y . z^z"

( "\frac {x+y+z} {3}" )(x+y+z)

=> ("\frac {x+y+z} {3}" )(x+y+z)"\ x^x . y^y . z^z"

("\frac {x²+y²+z²} {x+y+z}" )(x+y+z)